OpenStudy (anonymous):

The numerator of a fraction is 3 times the denominator. If 5 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 16/12

jhonyy9 (jhonyy9):

3d/d 3d-5 ----- = 16/12 so than 3d-5=16 so 3d=21 so d=7 so the numerator is 3d so is 21 5+d